# Homeostatic Adaptation of Optimal Population Codes under Metabolic Stress

### v1

@@ 4 Biophysical Simulation
We employ the biophysical simulator NEURON [12] to investigate the trade-offs between noise and energy. We simulate the response of a single-compartment neuron to an input spike, under a variety of settings. Specifically, we vary the three cellular parameters identified by [22] as changing under metabolic stress: resting potential $v_{rest}$, leak channel conductance $g_{leak}$, and synaptic conductance $g_{syn}$(see Fig.˜2a). Value ranges for $v_{rest}$ and $g_{leak}$ follow those of [22], and we vary $g_{syn}$ from 0 to 250 $\mu S/cm^{2}$. This range allows us to exclude cell parameter triplets that lead to a firing probability over .8 or under .2 in the deterministic setting, so that stochastic simulation captures meaningful variance. From $10,000$ simulated trials for each condition, we evaluate the mean and variance of the firing probability, denoted as $\mu_{F}$ and $\sigma_{F}^{2}$, as well as the energy expenditures associated with maintaining the resting state and signal activity (combining the cost of receiving and generating spikes), denoted as $\epsilon_{bg}$ and $\epsilon_{sig}$, respectively. See App.˜F for simulation parameters, App.˜G for details of energy accounting, and supplement for full simulation code.
We then derive higher level statistics from the single-spike characteristics (Fig. 2b). First, we compute the energy cost $\epsilon_{total}$ associated with each cell state. This cost combines rest state and signaling energy, weighted by an activity level $\kappa$ (see Sec.˜H.1). This factor is key for modeling visual neurons, which have orders-of-magnitude changes in spike rates along the visual pathway [9]. Then, we use the variance of spiking probability to characterize noise dispersion, details in the following section. Finally, we follow [22] to measure tuning curve widening $w$ by asserting a Gaussian tuning curve via rescaled synaptic conductance $g_{syn}$, details in Sec.˜H.2. Note that this procedure associates a unique value of $g_{syn}$ to each $(v_{rest},g_{leak})$ pair, reducing the degrees of freedom in our cell parameter space.
@@ 4.1 Noise-energy optimality
Figure 3: Simulation results of optimal adaptations under varying signal activity levels and intermediate outcomes. (a–c) Simulated results of total energy, and the mean and variance of firing probability, respectively. (d) The energy corresponding to a fixed firing probability, illustrated by the intersection in (a). (e) Fitted relationship between the mean and variance of firing probability for an example pair of $v_{rest}$ and $g_{leak}$, used to define the dispersion $\eta_{\kappa}(v_{rest},g_{leak})$. (f) Optimal adaptations obtained numerically by minimizing dispersion along each energy contour using the fitted energy and dispersion. (g) Optimal adaptations shown across different levels of signal activity $\kappa$. (h, i) The resulting density and dispersion as a function of optimal energy, respectively.
Using the noise $\sigma_{F}^{2}$ and energy $\epsilon_{total}$ associated with each cell state $(v_{rest},g_{leak},\kappa)$, we assert two hypotheses in turn to predict biophysical and information theoretic changes under metabolic stress, sketched in Fig. 2c and illustrated in detail in Fig. 3. First, we assert a strict homeostasis constraint to extract cell parameters corresponding to constant mean firing probability (white plane in Fig. 3a) and extract the corresponding energy costs (Fig. 3b, energy costs across the extracted plane shown in d). We note in Fig. 3c that firing rate variance saturates and then drops as $g_{syn}$ increases, as the probability of firing approaches 1. The variance is quadratic in mean firing probability, consistent with a Bernoulli distribution, but scaled by a factor that grows under increased metabolic stress. Specifically, we fit the scale of the firing variance parabola for the dispersion $\eta_{\kappa}$:
@@
Fig.˜3e illustrates this fitting for a single ($v_{rest},g_{leak}$) state. Although we fit the $\eta_{\kappa}$ based on firing probability (i.e., single spikes), it can be generalized to spike trains (see Sec.˜H.3).
@@
To numerically solve this minimization problem, we fit the $\eta_{\kappa}$ with a degree $8$ polynomial in $v_{rest}$ and $g_{leak}$, and intersect with a degree 1 model of $\epsilon_{total}$ respectively (see Sec.˜H.4). Each $(v_{rest},g_{leak})$ cell state along the optimal path $\delta_{opt}(\epsilon,\kappa)$ corresponds to both an energy cost and a noise level, providing the empirical trade-off between these terms, as well as a tuning curve width. Note that the shape of this curve varies with activity level $\kappa$, as shown in Fig. 3g. As $v_{rest}$ becomes less negative, signaling becomes cheaper but maintaining resting potential is more expensive, so only active cells deviate significantly from -75 mV. Additionally, changing the balance of background vs. signaling energy terms changes the slope of constant-energy-contours, which intersects with the nonlinear noise landscape to generate paths of different curvature (see App.˜I).
This simulation fixes the unknown terms $\alpha$ and $\eta_{\kappa}$ from our mathematical model as follows. First, we note that our optimal densities in Tab. 1 are all proportional to $E^{\frac{1}{\alpha}}$. Fig. 3h shows that, across activity levels, the optimal density has an affine relationship with $\epsilon_{opt}$, so that when we define $E$ in units of ATP/s as
@@
it follows immediately that
@@
See App.˜J for fits of $\vec{a}(\kappa)$ values from simulation, as well as $\vec{b}(\kappa)$, $\vec{c}(\kappa)$, and $\eta_{0}(\kappa)$ defined below.
The noise dispersion along the optimal adaptations shown in Fig.˜3i follows the intuitive trend that a decrease in energy leads to an increase in noise. Higher activity and tighter energy constraints lead to larger changes in noise levels. In general, the noise dispersion takes the form
$\eta_{\kappa}(\epsilon_{opt})~=$ $~\eta_{0}(\kappa)+\frac{b_{1}(\kappa)}{\epsilon_{opt}-b_{2}(\kappa)}$ (18)
$=$ $~\eta_{0}(\kappa)+\frac{c_{1}(\kappa)}{E(\kappa)-c_{2}(\kappa)},$ (19)
showing a direct trade-off between noise and energy in both physical units and the analytical energy budget $E$. Returning to our optimized solutions in Tab.˜1, we note that noise only indirectly affects the optimal density and gain, by shaping the optimal path and therefore setting $\alpha$. It also directly scales the Fisher information by a factor of $\eta_{\kappa}(E)^{-1}$, allowing empirically-derived $\eta_{\kappa}$ values to function as a plug-and-play term when estimating Fisher information for subsequent analyses, such as the evaluation of perceptual thresholds.
@@ 5 Consistency with data
Here, we will show that our method accurately predicts the tuning curve flattening effect seen in mouse cortex [22], while existing models incorrectly predict shortening or widening. We assume an infomax-optimal Gaussian tuning curve for a uniformly-distributed stimulus, and explore predicted changes under metabolic stress (see App.˜L for details). Mouse data in [22], illustrated in Fig.˜4a, shows a $32\%$ widening of the tuning curve and a statistically insignificant change in firing rate, which we model as an exact match in spike count. In Fig.˜4b, our model reiterates the 32% widening and $29\%$ reduction in ATP consumption reported in [22]. To accomplish both simultaneously, we fit $a_{1}$ and $a_{2}$ from eq.˜16 to satisfy $a_{2}/(a_{1}\epsilon_{ctr})=0.19625$, where $\epsilon_{ctr}$ denotes the energy expenditure in the control condition. In the infomax case with a uniform prior, we achieve a perfect tuning curve match because our approximation of homeostasis is exact. See App.˜L for homeostatic errors in other settings.
We compare our method to existing, non-homeostatic frameworks [5, 30]. In Fig.˜4c, [5] constrains the mean firing rate directly, with a fixed tuning curve width derived from the (fixed) population size. Because the tuning curve cannot match the widening observed in data, we reduce the firing rate to match the tuning curve peak instead. This requires a direct $24\%$ reduction in the firing rate, violating homeostasis by -24% under our uniform prior assumption. In Fig.˜4d, [30] maintains a constant max firing rate while reducing coding capacity, with a $24\%$ reduction in coding capacity required to widen by 32%, resulting in a homeostatic violation of 32%.
Figure 4: Comparison of optimal tuning curves in control versus food-restricted mouse neocortex (L2/3). (a-d) The figure illustrates how an example tuning curve in an optimal population adapts to a tightening of energy-related constraints. Under metabolic stress, real tuning curves flatten (a, based on [22]). Our model (b, purple) predicts this flattening, while existing models predict either shortening (c, blue [5]) or widening (d, red [30]). (e) We compute the change in mean firing rate for each method under a uniform prior. Only our model exhibits firing rate homeostasis.
@@ 6 Discussion
Our work provides an analytical and simulation-grounded framework that describes how metabolically-stressed neurons can optimally change to produce noisier, cheaper spike trains rather than reducing their firing rates, consistent with recent empirical observations [22]. This perhaps-counterintuitive firing rate homeostasis under metabolic stress is the key to correctly modeling how real tuning curves change. The standard analytical approach of limiting spike rates can predict optimal populations under a single metabolic state, but fails to capture how codes adapt to changing conditions. Changes in energy availability are a fact of life for all sensing organisms, and a key factor in evolutionary fitness. Indeed, there is good evidence that neural circuits maintain spike rate homeostasis across a range of contexts, perhaps as an essential part of keeping activity in neural circuits balanced [11, 34].
We have characterized the noise-energy trade-off in a simulation that connects a tractable and generalizing mathematical framework directly to ATP use. It also makes specific, testable predictions about how basic biophysical properties of cells should change under metabolic stress, factoring in activity level changes across the visual pathway. Specifically, the optimal paths in Fig.˜3g predict specific relationships between resting potential and leak conductance for neurons of different spike rates as they face metabolic pressure. These properties are readily measured in patch-clamp experiments, and future work will perform these experiments in matched types of neurons from control and food-restricted animals.
@@
Tractable, grounded models of metabolic impacts on neural codes are needed in order to ask new kinds of optimization questions, interpret existing data, and incorporate accelerating progress in our understanding of metabolism. The new optimization questions we can ask include: how can population codes be optimized to handle a range of energy budgets? Does an animal’s lifetime or early life access to calories impact their optimal codes? What can we tell about the needs and priorities of an animal from how their neural codes change under metabolic stress? In short, we can consider optimality in terms of long-term survival strategies rather than at snapshot states. The answer to these questions may lead to reinterpretations of existing data that has been used as evidence of particular optimization functions, as in [18, 7]. And while [22] describes changes that occur on the timescale of weeks, it is possible that some metabolic stress responses occur more quickly, with implications for the interpretation of neural data from food-deprived animals, common in behavioral literature. Though our model is based heavily on data from [22], a rapidly emerging understanding of the importance of metabolic and homeostatic factors in visual system function and nervous system health (e.g., [28, 17, 4, 19, 13, 25]) suggests that more data in this area may be available soon.
Acknowledgments and Disclosure of Funding
We thank Mark Agrios for his valuable assistance and insightful discussions on the simulations during the early stages of this project. Yi-Chun Hung gratefully acknowledges support from the Taiwan–Northwestern Doctoral Scholarship.
@@ Appendix A Approximate Fisher Information
The first assumption is commonly adopted in the previous literature [5, 30] and the second one is implicitly used in [5]. Based on these assumptions, we can approximate the total Fisher information $FI(s)$ as:
@@
The approximation relies on the assumptions. In our analysis, we further assume that the constant $I_{\text{conv}}$ is equal to $1$, as it does not influence the analytical solution of the proposed framework. For the self-consistency, we also derive the first equality in App.˜B.
@@ Appendix C Approximate Homeostasis
Followed by the approximation in Prop.˜C.1, we further relax the approximation to the population level and assume the normalized $\hat{h}(x)$. Then, we obtain the approximate homeostasis constraint:
@@ D.2 Discrimax
where we denote the integral in 42 as a scalar $A_{disc}$. Note that $A_{disc}$ is constant with respect to $s$ once the integral has been performed, but will vary if $p(s)$, $R(s)$, or $\alpha$ changes. Because this term will not change the shape of any predicted distributions (e.g., gain, density, or Fisher information), we suppress these dependencies in our notation fo convenience.

### v2

@@ 4 Biophysical Simulation
We employ the biophysical simulator NEURON (hines2022neuron) to investigate the trade-offs between noise and energy. We simulate a single-compartment neuron based on a stochastic Hodgkin-Huxley model in response to an input spike, under a variety of settings. Specifically, we vary the three cellular parameters identified by padamsey2022neocortex as changing under metabolic stress: resting potential $v_{rest}$, leak channel conductance $g_{leak}$, and synaptic conductance $g_{syn}$. See Fig. 2 for an overview of the simulation pipeline, App. F for an intuitive explanation of each parameter’s effect as well as a full list of simulation parameters, App. G for details of energy accounting, and the supplement for full simulation code.
Value ranges for $v_{rest}$ and $g_{leak}$ follow those of padamsey2022neocortex, and we vary $g_{syn}$ from 0 to 250 $\mu S/cm^{2}$. This range allows us to exclude cell parameter triplets that lead to mean spike count over .8 or under .2 total during the 2 second simulation duration in the deterministic setting, so that stochastic simulation captures meaningful variance. From $10,000$ simulated trials for each condition, we evaluate the mean and variance of the spike count, denoted as $\mu_{F}$ and $\sigma_{F}^{2}$, as well as the energy expenditures associated with maintaining the resting state and signal activity (combining the cost of receiving and generating spikes), denoted as $\epsilon_{bg}$ and $\epsilon_{sig}$, respectively.
We then derive higher level statistics from the single-spike characteristics. First, we compute the energy cost $\epsilon_{total}$ associated with each cell state. This cost combines rest state and signaling energy, weighted by an activity level $\kappa$ (see Sec. H.1). This factor is key for modeling visual neurons, which have orders-of-magnitude changes in spike rates along the visual pathway (goris2014partitioning). Then, we use the variance of spike count to characterize noise dispersion, details in the following section. Finally, we follow padamsey2022neocortex to measure tuning curve widening $w$ by asserting a Gaussian tuning curve via rescaled synaptic conductance $g_{syn}$, details in Sec. H.2. Note that this procedure associates a unique value of $g_{syn}$ to each $(v_{rest},g_{leak})$ pair, reducing the degrees of freedom in our cell parameter space.
@@ 4.1 Noise-energy optimality
Figure 3: Simulation results of optimal adaptations under varying signal activity levels and intermediate outcomes. (a–c) Simulated results of total energy, and the mean and variance of spike count, respectively. (d) The energy corresponding to a fixed spike count, illustrated by the intersection in (a). (e) Fitted relationship between the mean and variance of spike count for an example pair of $v_{rest}$ and $g_{leak}$, used to define the dispersion $\eta_{\kappa}(v_{rest},g_{leak})$. (f) Optimal adaptations obtained numerically by minimizing dispersion along each energy contour using the fitted energy and dispersion. (g) Optimal adaptations shown across different levels of signal activity $\kappa$. (h, i) The resulting density and dispersion as a function of optimal energy, respectively.
Using the noise $\sigma_{F}^{2}$ and energy $\epsilon_{total}$ associated with each cell state $(v_{rest},g_{leak},\kappa)$, we assert two hypotheses in turn to predict biophysical and information theoretic changes under metabolic stress, see Fig. 3. First, we assert a strict homeostasis constraint to extract cell parameters corresponding to constant mean spike count (white plane in Fig. 3a) and extract the corresponding energy costs (Fig. 3b, energy costs across the extracted plane shown in d). We note in Fig. 3c that firing count variance saturates and then drops as $g_{syn}$ increases, as the probability of firing approaches 1. The variance is quadratic in mean spike count, consistent with a Bernoulli distribution, but scaled by a factor that grows under increased metabolic stress. Specifically, we fit the scale of the firing variance parabola for the dispersion $\eta_{\kappa}$:
@@
Fig. 3e illustrates this fitting for a single ($v_{rest},g_{leak}$) state. Although we fit the $\eta_{\kappa}$ based on mean and variance of spike count under single spikes, it can be generalized to spike trains (see Sec. H.3).
@@
To numerically solve this minimization problem, we fit $\eta_{\kappa}$ using a 17-parameter model and intersect it with a linear fit of $\epsilon_{total}$ (see Sec. H.4 for fitting details). Each $(v_{rest},g_{leak})$ cell state along the optimal path $\delta_{opt}(\epsilon,\kappa)$ corresponds to both an energy cost and a noise level, providing the empirical trade-off between these terms, as well as a tuning curve width. Note that the shape of this curve varies with activity level $\kappa$, as shown in Fig. 3g. As $v_{rest}$ becomes less negative, signaling becomes cheaper but maintaining resting potential is more expensive, so only active cells deviate significantly from -75 mV. Additionally, changing the balance of background vs. signaling energy terms changes the slope of constant-energy-contours, which intersects with the nonlinear noise landscape to generate paths of different curvature (see App. I).
This simulation fixes the unknown terms $\alpha$ and $\eta_{\kappa}$ from our mathematical model as follows. First, we note that our optimal densities in Tab. 1 are all proportional to $E^{\frac{1}{\alpha}}$. When we define $E$ in units of ATP/s as
@@
we get the optimal density scales as $d(s)\propto E^{\frac{1}{\alpha}}=\left(a_{1}(\kappa)\epsilon_{opt}+a_{2}(\kappa)\right)^{\frac{1}{\alpha}}$. Since Fig. 3h shows that, across activity levels, the optimal density has an affine relationship with $\epsilon_{opt}$, it follows immediately that
@@
See App. J for fits of $\vec{a}(\kappa)$ values from simulation, as well as $\vec{b}(\kappa)$, $\vec{c}(\kappa)$, and $\eta_{0}(\kappa)$ defined below.
The noise dispersion along the optimal adaptations shown in Fig. 3i follows the intuitive trend that a decrease in energy leads to an increase in noise. Higher activity and tighter energy constraints lead to larger changes in noise levels. In general, the noise dispersion takes the form
$\eta_{\kappa}(\epsilon_{opt})~=$ $~\eta_{0}(\kappa)+\frac{b_{1}(\kappa)}{\epsilon_{opt}-b_{2}(\kappa)}=\eta_{0}(\kappa)+\frac{c_{1}(\kappa)}{E(\kappa)-c_{2}(\kappa)},$ (18)
showing a direct trade-off between noise and energy in both physical units and the analytical energy budget $E$. Returning to our optimized solutions in Tab. 1, we note that noise only indirectly affects the optimal density and gain, by shaping the optimal path and therefore setting $\alpha$. It also directly scales the Fisher information by a factor of $\eta_{\kappa}(E)^{-1}$, allowing empirically-derived $\eta_{\kappa}$ values to function as a plug-and-play term when estimating Fisher information for subsequent analyses, such as the evaluation of perceptual thresholds.
@@ 5 Consistency with data
Here, we will show that our method accurately predicts the tuning curve flattening effect seen in mouse cortex (padamsey2022neocortex), while existing models incorrectly predict shortening or widening. We assume an infomax-optimal Gaussian tuning curve for a uniformly-distributed stimulus, and explore predicted changes under metabolic stress (see App. L for details). Mouse data in padamsey2022neocortex, illustrated in Fig. 4a, shows a $32\%$ widening of the tuning curve and a statistically insignificant change in firing rate, which we model as an exact match in spike count. In Fig. 4b, our model reiterates the 32% widening and $29\%$ reduction in ATP consumption reported in padamsey2022neocortex. To accomplish both simultaneously, we fit $a_{1}$ and $a_{2}$ from eq. 16 to satisfy $a_{2}/(a_{1}\epsilon_{ctr})=0.19625$, where $\epsilon_{ctr}$ denotes the energy expenditure in the control condition. In the infomax case with a uniform prior, we achieve a perfect tuning curve match because our approximation of homeostasis is exact. See App. L for homeostatic errors in other settings.
We compare our method to existing, non-homeostatic frameworks (ganguli2010implicit; wang2016efficient). In Fig. 4c, ganguli2010implicit constrains the mean firing rate directly, with a fixed tuning curve width derived from the (fixed) population size. Because the tuning curve cannot match the widening observed in data, we reduce the firing rate to match the tuning curve peak instead. This requires a direct $24\%$ reduction in the firing rate, violating homeostasis by -24% under our uniform prior assumption. In Fig. 4d, wang2016efficient maintains a constant max firing rate while reducing coding capacity, with a $24\%$ reduction in coding capacity required to widen by 32%, resulting in a homeostatic violation of 32%. See App. M for a discussion of model likelihoods.
Figure 4: Comparison of optimal tuning curves in control versus food-restricted mouse neocortex (L2/3). (a-d) The figure illustrates how an example tuning curve in an optimal population adapts to a tightening of energy-related constraints. Under metabolic stress, real tuning curves flatten (a, based on padamsey2022neocortex). Our model (b, purple) predicts this flattening, while existing models predict either shortening (c, blue ganguli2010implicit) or widening (d, red wang2016efficient). (e) We compute the change in mean firing rate for each method under a uniform prior. Only our model exhibits firing rate homeostasis.
@@ 6 Discussion
Changes in energy availability are a fact of life for all sensing organisms, and a key factor in evolutionary fitness. Indeed, there is good evidence that neural circuits maintain spike rate homeostasis across a range of contexts, perhaps as an essential part of keeping activity in neural circuits balanced (Hengen2013; Wu2020).
We have characterized the noise-energy trade-off in a simulation that connects a tractable and generalizing mathematical framework directly to ATP use. It also makes specific, testable predictions about how basic biophysical properties of cells should change under metabolic stress, factoring in activity level changes across the visual pathway. Specifically, the optimal paths in Fig. 3g predict specific relationships between resting potential and leak conductance for neurons of different spike rates as they face metabolic pressure. These properties are readily measured in patch-clamp experiments, and future work will perform these experiments in matched types of neurons from control and food-restricted animals.
In summary, we analytically derive and predict that both density (i.e., $1/\text{width}$) and gain scale on the order of $E^{1/\alpha}$, while the Fisher information scales on the order of $\eta_{\kappa}(E)^{-1}\,E^{3/\alpha}$ (see Sec. 3). Additionally, we demonstrate that our model encompasses the competing models (ganguli2010implicit; wang2016efficient) as special cases. As a result, it preserves their explanatory power and, in particular, matches the ability of ganguli2016neural to produce signatures of optimality that have been found in brain areas for five measured attributes (three visual and two auditory). In Sec. 4, we use a stochastic Hodgkin–Huxley-like model to calibrate the parameters, yielding $\alpha=1$ and $\eta_{\kappa}(E)=\eta_{0}+\frac{c_{1}}{E-c_{2}}$. Finally, in Sec. 5, we show that the analytical solution with the calibrated $\alpha$ and $\eta_{\kappa}$ accurately predicts the tuning-curve flattening effect reported in padamsey2022neocortex.
@@
Tractable, grounded models of metabolic impacts on neural codes are needed in order to ask new kinds of optimization questions, interpret existing data, and incorporate accelerating progress in our understanding of metabolism. The new optimization questions we can ask include: how can population codes be optimized to handle a range of energy budgets? Does an animal’s lifetime or early life access to calories impact their neural codes? What can we tell about the needs and priorities of an animal from how their codes change under metabolic stress? In short, we can consider optimality in terms of long-term survival strategies rather than at snapshot states. The answer to these questions may lead to reinterpretations of existing data that has been used as evidence of particular optimization functions, as in manning2024transformations; ganguli2016neural. And while padamsey2022neocortex describes changes that occur on the timescale of weeks, it is possible that some metabolic stress responses occur more quickly, with implications for the interpretation of neural data from food-deprived animals, common in behavioral literature. Though our model is based heavily on data from padamsey2022neocortex, a rapidly emerging understanding of the importance of metabolic and homeostatic factors in visual system function and nervous system health (walls2025fueling; lin2025hypothalamus; etchegaray2011role; meng2025glucose; kim2025brain; sian2024metabolic) suggests that more data in this area may be available soon.
Acknowledgments
This work was supported by Taiwan-Northwestern Doctoral Scholarship. We would like to acknowledge Mark Agrios and the NSF-Simons National Institute for Theory and Mathematics in Biology (NITMB) for valuable discussions.
@@ Appendix A Approximate Fisher Information
The first assumption is commonly adopted in the previous literature (ganguli2010implicit; wang2016efficient) and the second one is implicitly used in ganguli2010implicit. Based on these assumptions, we can approximate the total Fisher information $FI(s)$ as:
@@
The approximation relies on the assumptions. In our analysis, we further assume that the constant $I_{\text{conv}}$ is equal to $1$, as it does not influence the analytical solution of the proposed framework. For the self-consistency, we also derive the first equality in App. B.
@@ Appendix C Approximate Homeostasis
Followed by the approximation in Prop. C.1, we further relax the approximation to the population level and assume the normalized $\hat{h}(x)$. Then, we obtain the approximate homeostasis constraint:
@@ D.2 Discrimax
where we denote the integral in eq. 41 as a scalar $A_{disc}$. Note that $A_{disc}$ is constant with respect to $s$ once the integral has been performed, but will vary if $p(s)$, $R(s)$, or $\alpha$ changes. Because this term will not change the shape of any predicted distributions (e.g., gain, density, or Fisher information), we suppress these dependencies in our notation fo convenience.